# MLlm-DR: Towards Explainable Depression Recognition with MultiModal Large Language Models

### v1

@@ 4.1. Datasets
We conduct experiments on two interview-based multimodal depression datasets (CMDC (CMDC, ) and E-DAIC-WOZ (DAIC, )) to evaluate the proposed method. Both datasets provide raw speech and text data, as well as scale-based scoring results. A score $\geq$ 10 indicates that the participant is experiencing severe depression. Additionally, for patient privacy protection, both datasets only provide the extracted visual features. CMDC: A Chinese dataset containing 78 samples (26 depressed patients and 52 healthy individuals), with PHQ-9 used as the corresponding questionnaire. Some subjects did not complete the video recording, resulting in missing visual features. E-DAIC-WOZ: An extended version of the DAIC-WOZ dataset, including 163 training samples, 56 validation samples, and 56 test samples, with PHQ-8 as the corresponding questionnaire. Our model is designed to assess each aspect independently and then aggregate the results to predict the overall depression score. Therefore, we treat each aspect-level response as an individual training instance, resulting in a total of $78\times 9=702$ training samples in CMDC and $163\times 8=1304$ in E-DAIC-WOZ.
@@ 4.2. Implementation Details
The method use LLaMA-3-8B (llama, ) as the base model with fine-tuning and functional extensions, and GPT-4o (gpt4, ) to construct the training datasets. The speech features are deep representations of 768-dim, while the visual features are deep representations of 709-dim or 2048-dim. For missing sequence features, zero padding is applied. LQ-former uses 32 query vectors, each with a dimension of 768-dim, a hidden layer dimension of 1024-dim, and 4 network layers, with a dropout rate set to 0.3. The projection layer employs a two-layer fully connected network with a hidden layer dimension of 1024-dim and an output dimension of 4096-dim. MAA uses 8 fully connected networks to learn the sequence weights for each attention head, followed by a two-layer fully connected network for regression prediction. We fine-tune the query and value projection matrices ($W_{q}$ and $W_{v}$) using LoRA by setting $r=16$, $\alpha=32$ and dropout = 0.1. The learning rate is set to 0.00001, and the model is trained for 10 epochs. All experiments are conducted using the PyTorch deep learning framework, and the training is performed on two A800 GPUs.
@@ 5.1. Baseline Methods
In our experiments, we introduce the following baseline methods for comparison: 1) CubeMLP(cubemlp, ): Utilizes multilayer perceptrons (MLPs) to perform feature fusion across multiple dimensions, enabling comprehensive interactions between modalities. 2) MulT(mult, ): A cross-attention-based multimodal fusion method designed to model interactions across different modalities. 3) MMFF(MMFF, ): Extracting multi-order factors from different modalities and their combinations, providing richer representational capacity for multimodal learning. 4) IIFDD(iifdd, ): Integrating intra-modality and inter-modality feature fusion frameworks for multimodal depression recognition. 5) HiQuE(HiQuE, ): Enhancing depression diagnosis accuracy by analyzing the importance weights of different questions within each modality. 6) LLaMA-3-8B(llama, ): Directly using the LLaMA-3-8B model to analyze dialogue content and generate evaluation results. 7) LoRA(lora, ): Fine-tuning the LLaMA-3-8B with the LoRA method using the training dataset. 8) P+W+A(Harnessing, ): Using prompts to guide LLMs in extracting textual features from transcripts, which are then fused with other modality features (AU, pose, gaze) for multimodal fusion.
@@ 5.2. Comparative Experimental
Table 2. Comparison of the performance of different methods on the CMDC and E-DAIC-WOZ datasets (i.e., {T} indicates that only the text modality was used, {A+T} indicates that both audio and text modalities were used, while others indicate that all modalities were used, including text, audio, and visual data. indicates methods that are based on LLMs).
Data Model CCC $\uparrow$ Pre $\uparrow$ Rec $\uparrow$ F1 $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$
CMDC CubeMLP \ 0.38 0.8 0.51 \ \
MulT {A+T} \ 0.72 \ \ 4.59 3.66
MMFF \ 0.83 \ \ 4.29 3.19
IIFDD \ 0.95 0.93 0.93 \ \
LLaMA-3-8B {T} 0.37 0.35 1 0.52 10.19 9.36
@@
Ours 0.91 1 1 1 3.10 2.61
E-DAIC- WOZ CubeMLP 0.58 \ \ \ \ 4.37
MulT \ 0.64 0.65 0.64 \ \
MMFF 0.67 \ \ \ 4.91 3.98
HiQuE \ 0.71 0.70 0.70 \ \
LLaMA-3-8B {T} 0.54 0.56 0.85 0.68 5.55 4.23
@@
We compared our proposed method with the latest in inroaches on the CMDC and E-DAIC-WOZ interview-based depression datasets, as shown in Table  2. The results demonstrate that our method outperforms all existing schemes, achieving state-of-the-art performance. Specifically, on the CMDC dataset, our method achieved an exceptional 100% in Precision, Recall, and F1 Score. On the E-DAIC-WOZ dataset, it also significantly outperformed other methods across all metrics. This superior performance is attributed to the LLMs’s powerful text understanding capabilities, which enable it to analyze the participant’s psychological state in specific aspects based on multi-turn dialogue content. Compared to the P+W+A in inroach (using LLMs), our method retains its advantage by integrating multimodal information and LLMs within a unified framework, reducing potential information loss in staged processing pipelines. Directly using LLaMA-3-8B for prediction does not perform well, primarily due to two factors:1) the limited logical reasoning capabilities of smaller LLMs, which make it challenging to provide accurate evaluations based on dialogue content; 2) the inherent limitations of language models in regression prediction tasks, stemming from a lack of domain-specific knowledge and sufficient labeled sample guidance. However, these issues were significantly mitigated through fine-tuning on training datasets, leading to substantial performance improvements.
@@ 5.3. Ablation Study
To verify the effectiveness of the LQ-former module and the multi-task learning strategy, we conduct extensive ablation experiments, as shown in Tables  3 and  4. The experimental settings include: 1) w/o Joint: Uses the LQ-former module to integrate audio and visual information and fine-tunes the LLMs, without multitask learning. 2) w/o LQ: Excludes the LQ-former module, not integrating audio and visual information, but fine-tunes with joint optimization. 3) w/o LQ-A: Excludes the audio LQ-former module but retains the visual LQ-former module, fine-tuning with joint optimization. 4) w/o LQ-V: Excludes the visual LQ-former module but retains the audio LQ-former module, fine-tuning with joint optimization. Experimental results demonstrate that the LQ-former module and the multi-task learning strategy can significantly enhance model performance, and removing either one would result in a performance decline. The performance improvement of w/o Joint is due to the pre-trained LQ-former module, which extracts depression-related features from audio and visual data, enhancing the LLM’s ability to process cross-modal information. The performance improvement of w/o LQ benefits from the mutual promotion between multiple tasks. The introduction of a regression prediction task enables the LLMs to generate scores that more closely align with the true values, even in the absence of sufficient prompt samples. Additionally, removing either audio or visual data in the LQ-former module also leads to a slight decline in model performance.
Table 3. Result of ablation study on the CMDC dataset.
Data Model CCC Pre Rec F1 RMSE MAE
CMDC w/o Joint 0.89 1 1 1 3.58 3.08
w/o LQ 0.87 1 1 1 4.18 3.61
w/o LQ-A 0.87 1 1 1 3.99 3.55
w/o LQ-V 0.90 1 1 1 3.52 2.90
Ours 0.91 1 1 1 3.10 2.61
Table 4. Result of ablation study on the E-DAIC-WOZ dataset.
Data Model CCC Pre Rec F1 RMSE MAE
@@ 5.4. Analysis of LQ-former
To validate the effectiveness of the features extracted by the LQ-former module from visual and speech data, we report the pre-training results of the LQ-former module, as shown in Tables  5 and  6. The feature representations extracted by LQ-former are fed into a a froze LLMs to generate depression scores, in the format of ”Score: 0”. The quality of the generated results reflects the LQ-former module’s performance in two aspects: first, its ability to extract depression-related features from speech and visual data; second, whether these features can be effectively understood by the LLMs. We did not have the model output the corresponding evaluation rationales, as generating fine-grained explanations solely based on speech and visual information is challenging. The main reason is that speech and visual data lack large-scale, fine-grained emotional labels, which prevents the LLMs from generating detailed and accurate explanations as it would with text data.
@@
Table 5. Result of LQ-former study on the CMDC dataset.
Data Model CCC Pre Rec F1 RMSE MAE
CMDC LQ-A 0.81 0.94 1 0.97 4.01 3.26
@@
LQ 0.85 1 1 1 3.79 3.06
Table 6. Result of LQ-former study on the E-DAIC-WOZ dataset.
Data Model CCC Pre Rec F1 RMSE MAE
@@ 5.5. Analysis of Evaluation Rationales
To further evaluate the interpretability and reliability of the model’s outputs, we conducted a human evaluation involving clinical experts. A total of 100 samples were randomly selected from the test set. Experts were asked to independently score each sample and evaluate the model-generated evaluation rationales corresponding to those samples. Each rationale was rated on a 3-point scale: 3 — Fully agree (the expert would have made the same assessment), 2 — Reasonable (different perspective but similar conclusion), and 1 — Disagree (the rationale was not acceptable). We report four evaluation metrics in Table  7: I) The proportion of model outputs not conforming to the required instruction format; II) Result of expert ratings (RMSE/MAE); III) agreement rate between model predictions and expert ratings; IV) Proportional distribution of expert ratings on the quality of model-generated evaluation rationales (3/2/1). From the results, we draw the following observations: 1) LLaMA-3-8B frequently produced improperly formatted outputs, impairing scoring accuracy, while LoRA fine-tuning effectively mitigated this issue and enhanced model performance. 2) Our method achieved 87% and 73% consistency with expert scores on CMDC and E-DAIC-WOZ respectively, demonstrating strong practical value. 3) Expert ratings also showed limitations, even falling slightly behind our method on E-DAIC-WOZ, reflecting the inherent difficulty of depression assessment and the subtlety of depressive cues that often require additional context to interpret. 4) 77% and 64% of the model’s reasoning was rated as “fully agree” by experts on CMDC and E-DAIC-WOZ respectively, highlighting the interpretability of our approach.
@@
IV 0.77/0.14/0.09 0.64/0.19/0.17
@@ 5.6. Case Study
1) LLaMA-3-8B method showed clear deficiencies in score prediction. For instance, although the model recognized that the subject had 8 hours of sleep, it still gave a score of 2 (poor), which did not align with the actual situation. 2) LoRA method fine-tuned model showed improvements, providing a ”mild” reasoning result, but the gap between the predicted scores and actual values remained significant. This limitation is primarily due to the lack of real labels during training, leading to deviations between predicted and actual scores. 3) Our method demonstrated a clear advantage in capturing subtle cues in the dialogue that are difficult to detect, enabling accurate predictions. This advantage is attributed to the integration of multi-modal and multi-task learning modules, which provide the model with latent information, thereby enhancing the accuracy of score predictions.

### v2

@@ 4.1. Datasets
We conduct experiments on two interview-based multimodal depression datasets (CMDC (Zou et al., 2022) and E-DAIC-WOZ (Gratch et al., 2014)) to evaluate the proposed method. Both datasets provide raw speech and text data, as well as scale-based scoring results. A score $\geq$ 10 indicates that the participant is experiencing severe depression. Additionally, for patient privacy protection, both datasets only provide the extracted visual features. CMDC: A Chinese dataset containing 78 samples (26 depressed patients and 52 healthy individuals), with PHQ-9 used as the corresponding questionnaire. Some subjects did not complete the video recording, resulting in missing visual features. E-DAIC-WOZ: An extended version of the DAIC-WOZ dataset, including 163 training samples, 56 validation samples, and 56 test samples, with PHQ-8 as the corresponding questionnaire.
Our model is designed to assess each aspect independently and then aggregate the results to predict the overall depression score. Therefore, we treat each aspect-level response as an individual training instance, resulting in a total of $78\times 9=702$ training samples in CMDC and $163\times 8=1304$ in E-DAIC-WOZ. For both datasets, we adopt subject-independent splits, ensuring that all samples from the same participant (e.g., different PHQ items) appear in only one split, thus avoiding information leakage.
@@ 4.2. Implementation Details
The method use LLaMA-3-8B (Touvron et al., 2023) as the base model with fine-tuning and functional extensions, and GPT-4o (Achiam et al., 2023) to construct the training datasets. The speech features are deep representations of 768-dim, while the visual features are deep representations of 709-dim or 2048-dim. For missing sequence features, zero padding is applied. LQ-former uses 32 query vectors, each with a dimension of 768-dim, a hidden layer dimension of 1024-dim, and 4 network layers, with a dropout rate set to 0.3. The projection layer employs a two-layer fully connected network with a hidden layer dimension of 1024-dim and an output dimension of 4096-dim. MAA uses 8 fully connected networks to learn the sequence weights for each attention head, followed by a two-layer fully connected network for regression prediction. We fine-tune the query and value projection matrices ($W_{q}$ and $W_{v}$) using LoRA by setting $r=16$, $\alpha=32$ and dropout = 0.1. The learning rate is set to 0.00001, and the model is trained for 10 epochs. Text generation uses nucleus sampling with do_sample=True, temperature=0.8, and top_k=10. All experiments are conducted using the PyTorch deep learning framework, and the training is performed on two A800 GPUs.
@@ 5.1. Baseline Methods
In our experiments, we introduce the following baseline methods for comparison: 1) CubeMLP(Sun et al., 2022a): Utilizes multilayer perceptrons (MLPs) to perform feature fusion across multiple dimensions, enabling comprehensive interactions between modalities. 2) MulT(Zadeh et al., 2019): A cross-attention-based multimodal fusion method designed to model interactions across different modalities. 3) MMFF(Yuan et al., 2024): Extracting multi-order factors from different modalities and their combinations, providing richer representational capacity for multimodal learning. 4) IIFDD(Chen et al., 2024): Integrating intra-modality and inter-modality feature fusion frameworks for multimodal depression recognition. 5) HiQuE(Jung et al., 2024): Enhancing depression diagnosis accuracy by analyzing the importance weights of different questions within each modality. 6) P+W+A(Sadeghi et al., 2024): Using prompts to guide LLMs in extracting textual features from transcripts, which are then fused with other modality features (AU, pose, gaze) for multimodal fusion. 7) MMCL (Wang et al., 2025): captures enhanced and complementary features from both modality-common and modality-specific representations, which are then fused into a unified representation for emotion prediction.
@@ 5.2. Comparative Experimental
We compared our proposed method with the latest in inroaches on the CMDC and E-DAIC-WOZ interview-based depression datasets, as shown in Table  2. The results demonstrate that our method outperforms all existing schemes, achieving state-of-the-art performance. Specifically, on the CMDC dataset, our method achieved an exceptional 100% in Precision, Recall, and F1 Score. On the E-DAIC-WOZ dataset, it also significantly outperformed other methods across all metrics, confirming the robustness and generalization ability of the proposed framework. This superior performance can be attributed to the LLM’s powerful contextual understanding capabilities, which enable it to capture fine-grained semantic and emotional cues from multi-turn dialogues, thereby allowing a more accurate analysis of the participant’s psychological state in specific aspects.
Table 2. Comparison of the performance of recent methods on the CMDC and E-DAIC-WOZ datasets, Bold numbers indicate the best performance.
Dataset Model CCC $\uparrow$ Pre $\uparrow$ Rec $\uparrow$ F1 $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$
CMDC MulT(2019) \ 0.72 \ \ 4.59 3.66
CubeMLP(2022) \ 0.38 0.8 0.51 \ \
MMFF(2022) \ 0.83 \ \ 4.29 3.19
IIFDD(2023) \ 0.95 0.93 0.93 \ \
MMCL(2025) \ 1.00 0.93 0.96 3.43 2.31
Ours 0.91 1.00 1.00 1.00 3.10 2.61
E-DAIC- WOZ MulT(2019) \ 0.64 0.65 0.64 \ \
CubeMLP(2022) 0.58 \ \ \ \ 4.37
MMFF(2022) 0.67 \ \ \ 4.91 3.98
HiQuE(2024) \ 0.71 0.70 0.70 \ \
P+W+A(2024) \ \ \ \ 4.66 3.86
MMCL(2025) \ 0.74 0.73 0.72 4.76 3.85
Ours 0.72 0.77 0.81 0.79 4.59 3.41
Compared with non-LLMs-based methods such as MulT (Zadeh et al., 2019), CubeMLP (Sun et al., 2022a), MMFF (Yuan et al., 2024), IIFDD (Chen et al., 2024), HiQuE (Jung et al., 2024), and MMCL (Wang et al., 2025), our model exhibits substantial advantages. These methods typically rely on small language models such as BERT (Devlin et al., 2019) for textual feature extraction, which limits their contextual awareness in multi-turn dialogues and consequently constrains their overall performance. Compared with LLMs-based approaches such as P+W+A (Sadeghi et al., 2024), our method still achieves superior overall results. P+W+A first uses an LLMs to extract textual representations and then performs fusion with acoustic and visual embeddings. In contrast, our method retains its advantage by integrating multimodal information and LLMs within a unified framework, reducing potential information loss in staged processing pipelines.
5.3. Effectiveness of the Training Dataset
To evaluate the effectiveness of the constructed training dataset, we conduct comparative experiments on two baseline methods, as shown in Table 3. 1) LLaMA-3-8B(Touvron et al., 2023): Directly using the LLaMA-3-8B model to analyze dialogue content and generate evaluation results. 2) LoRA(Hu et al., 2022): Fine-tuning the LLaMA-3-8B with the LoRA method using the training dataset.
Table 3. Comparison of the performance of different baselines on the CMDC and E-DAIC-WOZ datasets (i.e., {T} indicates that only the text modality was used).
Dataset Model CCC $\uparrow$ Pre $\uparrow$ Rec $\uparrow$ F1 $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$
CMDC LLaMA-3-8B {T} 0.37 0.35 1.00 0.52 10.19 9.36
@@ 5.3. Effectiveness of the Training Dataset
Ours 0.91 1.00 1.00 1.00 3.10 2.61
E-DAIC- WOZ LLaMA-3-8B {T} 0.54 0.56 0.85 0.68 5.55 4.23
@@
Directly using LLaMA-3-8B for prediction does not perform well, primarily due to three factors: 1) the limited logical reasoning capabilities of small-scale LLMs , which make it challenging to provide accurate evaluations based on dialogue content; 2) the model often fails to strictly follow instruction-defined scoring formats, leading to unusable outputs; 3) the inherent limitations of language models in regression prediction tasks, stemming from a lack of domain-specific knowledge and sufficient labeled sample guidance. However, these issues were significantly mitigated through LoRA-based fine-tuning on training datasets, leading to substantial performance improvements.
@@ 5.4. Ablation Study
To verify the effectiveness of the LQ-former module and the multi-task learning strategy, we conduct extensive ablation experiments, as shown in Table  4. The experimental settings include: 1) w/o Joint: Uses the LQ-former module to integrate audio and visual information and fine-tunes the LLMs, without multitask learning. 2) w/o LQ: Excludes the LQ-former module, not integrating audio and visual information, but fine-tunes with joint optimization. 3) w/o LQ-A: Excludes the audio LQ-former module but retains the visual LQ-former module, fine-tuning with joint optimization. 4) w/o LQ-V: Excludes the visual LQ-former module but retains the audio LQ-former module, fine-tuning with joint optimization.
Experimental results demonstrate that the LQ-former module and the multi-task learning strategy can significantly enhance model performance, and removing either one would result in a performance decline. The LQ-former is capable of extracting depression-related cues from acoustic and visual data, thereby providing the LLMs with more informative cross-modal representations. Meanwhile, the joint learning strategy introduces a regression prediction task that enables the LLMs to generate scores more closely aligned with the true values, even when prompt examples are insufficient. Additionally, removing either audio or visual data in the LQ-former module also leads to a slight decline in model performance.
Table 4. Result of ablation study on the CMDC and E-DAIC-WOZ datasets.
Dataset Model CCC $\uparrow$ Pre $\uparrow$ Rec $\uparrow$ F1 $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$
CMDC w/o Joint 0.89 1.00 1.00 1.00 3.58 3.08
w/o LQ 0.87 1.00 1.00 1.00 4.18 3.61
w/o LQ-A 0.87 1.00 1.00 1.00 3.99 3.55
w/o LQ-V 0.90 1.00 1.00 1.00 3.52 2.90
Ours 0.91 1.00 1.00 1.00 3.10 2.61
@@ 5.5. Analysis of LQ-former
To validate the effectiveness of the features extracted by the LQ-former module from visual and speech data, we report the pre-training results of the LQ-former module, as shown in Table  5. The feature representations extracted by LQ-former are fed into a a froze LLMs to generate depression scores, in the format of ”Score: 0”. The quality of the generated results reflects the LQ-former module’s performance in two aspects: first, its ability to extract depression-related features from speech and visual data; second, whether these features can be effectively understood by the LLMs. We did not have the model output the corresponding evaluation rationales, as generating fine-grained explanations solely based on speech and visual information is challenging. The main reason is that speech and visual data lack large-scale, fine-grained emotional labels, which prevents the LLMs from generating detailed and accurate explanations as it would with text data.
@@
Table 5. Result of LQ-former study on the CMDC and E-DAIC-WOZ datasets.
Dataset Model CCC $\uparrow$ Pre $\uparrow$ Rec $\uparrow$ F1 $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$
CMDC LQ-A 0.81 0.94 1.00 0.97 4.01 3.26
@@
LQ 0.85 1.00 1.00 1.00 3.79 3.06
@@ 5.6. Analysis of Evaluation Rationales
To further evaluate the interpretability and reliability of the model’s outputs, we conducted a human evaluation involving clinical experts. We randomly selected 100 samples from the test set and invited three assessors with mental-health–related professional backgrounds who had received formal training from physicians to perform independent reviews. The experts evaluated the corresponding model-generated rationales. Each rationale was rated on a 3-point scale: 3 — Fully agree (the expert would have made the same assessment), 2 — Reasonable (different perspective but similar conclusion), and 1 — Disagree (the rationale was not acceptable). The final per-sample rationale label was determined by majority vote; in the event of a tie, the median rating was used. We report four evaluation metrics in Table  6: I) The proportion of model outputs not conforming to the required instruction format; II) agreement rate between model predictions and expert ratings; III) Result of expert ratings (RMSE/MAE); IV) Proportional distribution of expert ratings on the quality of model-generated evaluation rationales (3/2/1).
@@
IV 77%/14%/9% 64%/19%/17%
From the results, we draw the following observations: 1) LLaMA-3-8B frequently produced improperly formatted outputs, impairing scoring accuracy, while LoRA fine-tuning effectively mitigated this issue and enhanced model performance. 2) Our method achieved 87% and 73% consistency with expert scores on CMDC and E-DAIC-WOZ respectively, demonstrating strong practical value. 3) Expert ratings also showed limitations, even falling slightly behind our method on E-DAIC-WOZ, reflecting the inherent difficulty of depression assessment and the subtlety of depressive cues that often require additional context to interpret. (4) On CMDC and E-DAIC-WOZ, 77% and 64% of the model’s rationales were rated “fully agree” by clinical experts, further corroborating the interpretability and clinical concordance of the proposed approach.
@@ 5.7. Case Study
The results reveal a progressive improvement across methods: 1) LLaMA-3-8B method showed clear deficiencies in score prediction. For instance, although the model recognized that the subject had 8 hours of sleep, it still gave a score of 2 (poor), which did not align with the actual situation. 2) LoRA method fine-tuned model showed improvements, providing a ”mild” reasoning result, but the gap between the predicted scores and actual values remained significant. This limitation is primarily due to the lack of real labels during training, leading to deviations between predicted and actual scores. 3) Our method demonstrated a clear advantage in capturing subtle cues in the dialogue that are difficult to detect, enabling accurate predictions. This advantage is attributed to the integration of multimodal and multi-task learning modules, which provide the model with latent information, thereby enhancing the accuracy of score predictions.
5.8. Efficiency Analysis
We further analyze the efficiency of our method. Specifically, we adopt a two-stage training strategy. In Stage 1, the LLMs is frozen and only the LQ-former is trained. The LQ-former consists of four Transformer decoder layers (hidden size = 1024) with 32 learnable queries and a two-layer projection module. The additional computational overhead in this stage primarily comes from multimodal feature encoding. In Stage 2, we introduce LoRA adaptation to the ($W_{q}$ and $W_{v}$) matrices of LLaMA-3-8B with ($r=16$) and ($\alpha=32$). The number of trainable parameters is on the order of several million, which is three orders of magnitude smaller than full-parameter fine-tuning ($\approx$8B parameters). Consequently, the overall training cost is only slightly higher than that of standard LoRA-based baselines, and during inference our method adds merely a single forward pass of the LQ-former, whose latency is negligible compared to the LLMs. A detailed comparison of trainable parameters, GPU memory, and training time is provided in Table 7.
Despite this minimal extra cost, our method yields substantial performance improvements over the LoRA baseline, as shown in Table 3. On E-DAIC-WOZ, CCC increases from 0.69 to 0.72, F1 from 0.77 to 0.79, and MAE decreases from 4.04 to 3.41; on CMDC, RMSE drops from 4.61 to 3.10, MAE from 4.08 to 2.61, and all three classification metrics reach 1.00. These results demonstrate that the proposed two-stage strategy offers a highly favorable performance–efficiency trade-off, delivering stable and substantial gains with only marginal additional training overhead.
Table 7. Comparison of training cost between different models on the CMDC dataset.
Dataset Model Trainable Params GPU Memory Training Time
CMDC MulT 20M 16G 1.5h
LLaMA-3–8B 8M 80G 2.5h
Ours 25M 80G 3.0h
5.9. Model Scalability and Efficiency
To analyze the impact of model size, we instantiated MLlm-DR with LLaMA-3 models of different parameter scales (1B, 3B, and 8B) while keeping all training settings and the LQ-former unchanged. As shown in Table 8, performance declines consistently as the model size decreases. The 3B model exhibits only a mild degradation relative to the 8B baseline, whereas the 1B model shows substantial drops across all metrics, with CCC and F1 decreasing by 25.3% and 61.1%, respectively. These findings indicate that an 8B-scale LLMs provides a favorable balance between reasoning capability and computational overhead for explainable depression recognition, while excessively small models are insufficient to capture the nuanced semantic and emotional cues present in clinical interviews.
Table 8. Performance of MLlm-DR with different scales LLM on the E-DAIC-WOZ dataset.
Dataset Model CCC $\uparrow$ Pre $\uparrow$ Rec $\uparrow$ F1 $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$
E-DAIC- WOZ LLaMA-3-1B 0.28 0.52 0.71 0.59 6.05 4.94
LLaMA-3-3B 0.65 0.76 0.54 0.63 4.70 3.59
LLaMA-3-8B 0.72 0.77 0.81 0.79 4.59 3.41
5.10. Cross-Corpus Generalization Study
To further verify the generalization capability of MLlm-DR, we conducted a cross-corpus experiment training on the Chinese CMDC dataset and evaluating on the English E-DAIC-WOZ dataset. This setting challenges the model not only on domain shift but also on linguistic diversity. Specifically, to bridge the linguistic gap, the text modality was translated to align with the target language during the evaluation. As shown in Table  9, our model demonstrates robust performance in this challenging setting, outperforming the zero-shot baseline of LLaMA-3-8B. This confirms that our method possesses strong generalization capabilities across different languages and data distributions, even when constrained by feature heterogeneity.
Table 9. Cross-corpus evaluation: training on CMDC and testing on E-DAIC-WOZ.
Model CCC $\uparrow$ Pre $\uparrow$ Rec $\uparrow$ F1 $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$
zero-shot 0.54 0.56 0.85 0.68 5.55 4.23
cross-corpus 0.69 0.65 0.71 0.68 5.04 4.04
5.11. Effect of Different Speech Features
To investigate how different acoustic representations affect the multimodal reasoning ability of MLlm-DR, we replaced the original HuBERT features with two alternative extractors, Mel-spectrograms(McFee et al., 2015) and WavLM(Chen et al., 2022), and conducted controlled experiments on the CMDC dataset. As shown in Table 10, Mel-spectrogram features lead to a marked degradation across all metrics, suggesting that low-level spectral energy information alone is insufficient to capture depression-related speech patterns. WavLM outperforms Mel and yields moderate gains in CCC and RMSE, indicating that its self-supervised acoustic representations are more effective for modeling speaker state; however, its overall performance still falls short of HuBERT. HuBERT, trained with a “cluster + masked prediction” objective, preserves higher-level speech representations that are tightly coupled with semantic content while still encoding paralinguistic cues. Therefore, HuBERT remains the most suitable backbone for speech feature extraction in our clinical interview setting.
Table 10. Performance of MLlm-DR with different acoustic feature on the CMDC dataset.
Dataset Model CCC $\uparrow$ Pre $\uparrow$ Rec $\uparrow$ F1 $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$
CMDC Mel 0.40 0.55 0.83 0.67 7.78 6.28
WavLM 0.59 0.71 0.83 0.79 6.12 4.83
HuBERT 0.81 0.94 1.00 0.97 4.01 3.26
5.12. Stability and Robustness Analysis
To further evaluate the robustness and stability of MLlm-DR, we conducted five independent runs on both the CMDC and E-DAIC-WOZ datasets using the same LLaMA-3.2-8B backbone and identical training configurations. As summarized in Table 11, the model maintains consistently high performance across runs, with only minor fluctuations in CCC, F1, RMSE, and MAE. On CMDC, the average CCC reaches 0.926 and F1 remains at 0.964, demonstrating strong stability on interview data. On E-DAIC-WOZ, despite the dataset’s open-ended conversational nature, the model still achieves an average CCC of 0.646 and F1 of 0.688, confirming its reliable generalization behavior. The results confirm that the performance gains of MLlm-DR are stable and statistically meaningful, indicating that they stem from the method itself rather than stochastic fluctuations.
Table 11. Results of five independent runs of MLlm-DR on the CMDC and E-DAIC-WOZ dataset.
Dataset Run CCC $\uparrow$ Pre $\uparrow$ Rec $\uparrow$ F1 $\uparrow$ RMSE $\downarrow$ MAE $\downarrow$
CMDC 1 0.93 1.00 0.83 0.91 3.59 2.83
2 0.93 1.00 0.83 0.91 3.63 2.94
3 0.92 1.00 1.00 1.00 3.77 3.22
4 0.94 1.00 1.00 1.00 3.55 2.89
5 0.91 1.00 1.00 1.00 3.10 2.61
average 0.93 1.00 0.93 0.96 3.52 2.89
E-DAIC- WOZ 1 0.64 0.80 0.67 0.73 4.77 3.58
2 0.60 0.71 0.63 0.67 5.11 3.69
3 0.67 0.72 0.54 0.62 4.66 3.46
4 0.60 0.76 0.54 0.63 5.02 3.70
5 0.72 0.77 0.81 0.79 4.59 3.41
average 0.65 0.75 0.64 0.69 4.83 3.57